OpenStudy (anonymous):

Tamara needs to buy motor oil to fill the 3 empty cylindrical barrels at her oil service center. Each barrel is 7 ft deep and has a radius of 4 ft. What is the volume of oil needed?

OpenStudy (austinl):

\(V=\pi r^2 h\) We need \(3V\).

OpenStudy (anonymous):

V=πr2h pi = 3.14 , r = 4 ft , h = 7 ft Volume of cylinder is pi*r^2*h V = 3.14*4^2*7 = 351.68 ft^3 For 3 cylinders, V = 3*351.68 =1055.04 ft^3

OpenStudy (anonymous):

thanks!!!!!

OpenStudy (anonymous):

welcome!!!